Circular No. 8583

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

CBAT@CFA.HARVARD.EDU (science)

URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/1984 R1 (SOLWIND)

Further to IAUC 8573, R. Kracht reports another comet found and measured by himself from SOLWIND images.

1984 UT  α_{2000}  δ_{2000}
Sept.15.245  11^{h}29.^{m}4  + 4^{o}14′
15.318  11 30.4  + 4 14
15.372  11 31.1  + 4 13
15.375  11 31.2  + 4 14
15.378  11 31.2  + 4 13
15.713  11 35.9  + 4 11
15.717  11 35.9  + 4 11
15.780  11 36.8  + 4 11
15.783  11 36.9  + 4 11
16.172  11 42.6  + 4 08
16.175  11 42.6  + 4 07
16.179  11 42.7  + 4 08
16.182  11 42.7  + 4 08
16.310  11 44.7  + 4 06
16.313  11 44.7  + 4 07
16.365  11 45.5  + 4 06

The following parabolic orbital elements are by B. G. Marsden. The comet seems not to belong to a known group.

T = 1984 Sept.17.42 TT  ω = 78.^{o}74
Ω = 152.84 }2000.0
q = 0.1051 AU  i = 36.92

SUPERNOVA 2005dh in MCG +04-37-55

Further to IAUC 8580, M. Moore and W. Li report the LOSS discovery of an apparent supernova on unfiltered KAIT images taken on Aug. 10.22 (at mag 17.1) and 11.23 UT (mag 16.9). SN 2005dh is located at α = 15^{h}54^{m}03^{s}.32, δ = +23^{o}08′09′′.9 (equinox 2000.0), which is 6′′.1 east and 16′′.4 north of the nucleus of MCG +04-37-55. A KAIT image taken on Aug. 1.20 showed nothing at this position (limiting mag ∼ 19.0).

2005 August 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green